$SPAD/input schaum29.input

Timothy Daly December 3, 2016

Contents

2

1 [?]:14.590 ∫ sinh ax cosh ax dx

∫ sinh ax cosh ax = \frac{sinh^{2} ax}{2a}

— * —

)set break resume
)spool schaum29.output
)set message test on
)set message auto off
)clear all
--S 1 of 81
aa:=integrate(sinh(a*x)*cosh(a*x),x)
--R
--R
--R                 2            2
--R        sinh(a x)  + cosh(a x)
--R   (1)  -----------------------
--R                   4a
--R                                         Type: Union(Expression(Integer),...)
--E
--S 2 of 81
bb:=sinh(a*x)^2/(2*a)
--R
--R                 2
--R        sinh(a x)
--R   (2)  ----------
--R            2a
--R                                                    Type: Expression(Integer)
--E
--S 3 of 81
cc:=aa-bb
--R
--R                   2            2
--R        - sinh(a x)  + cosh(a x)
--R   (3)  -------------------------
--R                    4a
--R                                                    Type: Expression(Integer)
--E
--S 4 of 81
sinhsqrrule:=rule(sinh(x)^2 == 1/2*cosh(2*x)-1/2)
--R
--R               2    cosh(2x) - 1
--R   (4)  sinh(x)  == ------------

--R                          2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 5 of 81
dd:=sinhsqrrule cc
--R
--R                                 2
--R        - cosh(2a x) + 2cosh(a x)  + 1
--R   (5)  ------------------------------
--R                      8a
--R                                                    Type: Expression(Integer)
--E
--S 6 of 81
coshsqrrule:=rule(cosh(x)^2 == 1/2*cosh(2*x)+1/2)
--R
--R               2    cosh(2x) + 1
--R   (6)  cosh(x)  == ------------
--R                          2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 7 of 81      14:590 Schaums and Axiom differ by a constant
ee:=coshsqrrule dd
--R
--R         1
--R   (7)  --
--R        4a
--R                                                    Type: Expression(Integer)
--E

———-

2 [?]:14.591 ∫ sinh px cosh qx dx

∫ sinh px cosh qx = \frac{cosh(p + q)x}{2(p + q)} + \frac{cosh(p − q)x}{2(p − q)}

— * —

)clear all
--S 8 of 81
aa:=integrate(sinh(p*x)*cosh(q*x),x)
--R
--R
--R        - q sinh(p x)sinh(q x) + p cosh(p x)cosh(q x)
--R   (1)  ---------------------------------------------

--R           2    2          2       2    2          2
--R         (q  - p )sinh(p x)  + (- q  + p )cosh(p x)
--R                                         Type: Union(Expression(Integer),...)
--E
--S 9 of 81
bb:=(cosh(p+q)*x)/(2*(p+q))+(cosh(p-q)*x)/(2*(p-q))
--R
--R        (q - p)x cosh(q + p) + (- q - p)x cosh(q - p)
--R   (2)  ---------------------------------------------
--R                            2     2
--R                          2q  - 2p
--R                                                    Type: Expression(Integer)
--E
--S 10 of 81     14:591 Axiom cannot simplify this expression
cc:=aa-bb
--R
--R   (3)
--R       - 2q sinh(p x)sinh(q x)
--R     +
--R                                                               2
--R       ((- q + p)x cosh(q + p) + (q + p)x cosh(q - p))sinh(p x)
--R     +
--R       2p cosh(p x)cosh(q x)
--R     +
--R                                                               2
--R       ((q - p)x cosh(q + p) + (- q - p)x cosh(q - p))cosh(p x)
--R  /
--R        2     2          2        2     2          2
--R     (2q  - 2p )sinh(p x)  + (- 2q  + 2p )cosh(p x)
--R                                                    Type: Expression(Integer)
--E

———-

3 [?]:14.592 ∫ sinh^{n} ax cosh ax dx

∫ sinh^{n} ax cosh ax = \frac{sinh^{n+1} ax}{(n + 1)a}

— * —

)clear all
--S 11 of 81
aa:=integrate(sinh(a*x)^n*cosh(a*x),x)
--R
--R

--R        - sinh(a x)sinh(n log(sinh(a x))) - sinh(a x)cosh(n log(sinh(a x)))
--R   (1)  -------------------------------------------------------------------
--R                                      2                       2
--R                    (a n + a)sinh(a x)  + (- a n - a)cosh(a x)
--R                                         Type: Union(Expression(Integer),...)
--E
--S 12 of 81
bb:=sinh(a*x)/((n+1)*a)
--R
--R        sinh(a x)
--R   (2)  ---------
--R         a n + a
--R                                                    Type: Expression(Integer)
--E
--S 13 of 81     14:592 Axiom cannot simplify this expression
cc:=aa-bb
--R
--R   (3)
--R       - sinh(a x)sinh(n log(sinh(a x))) - sinh(a x)cosh(n log(sinh(a x)))
--R     +
--R                  3            2
--R       - sinh(a x)  + cosh(a x) sinh(a x)
--R  /
--R                       2                       2
--R     (a n + a)sinh(a x)  + (- a n - a)cosh(a x)
--R                                                    Type: Expression(Integer)
--E

———-

4 [?]:14.593 ∫ cosh^{n} ax sinh ax dx

∫ cosh^{n} ax sinh ax = \frac{cosh^{n+1} ax}{(n + 1)a}

— * —

)clear all
--S 14 of 81
aa:=integrate(cosh(a*x)^n*sinh(a*x),x)
--R
--R
--R        - cosh(a x)sinh(n log(cosh(a x))) - cosh(a x)cosh(n log(cosh(a x)))
--R   (1)  -------------------------------------------------------------------
--R                                      2                       2
--R                    (a n + a)sinh(a x)  + (- a n - a)cosh(a x)

--R                                         Type: Union(Expression(Integer),...)
--E
--S 15 of 81
bb:=cosh(a*x)^(n+1)/((n+1)*a)
--R
--R                 n + 1
--R        cosh(a x)
--R   (2)  --------------
--R            a n + a
--R                                                    Type: Expression(Integer)
--E
--S 16 of 81     14:593 Axiom cannot simplify this expression
cc:=aa-bb
--R
--R   (3)
--R       - cosh(a x)sinh(n log(cosh(a x))) - cosh(a x)cosh(n log(cosh(a x)))
--R     +
--R                   2            2          n + 1
--R       (- sinh(a x)  + cosh(a x) )cosh(a x)
--R  /
--R                       2                       2
--R     (a n + a)sinh(a x)  + (- a n - a)cosh(a x)
--R                                                    Type: Expression(Integer)
--E

———-

5 [?]:14.594 ∫ sinh^{2} ax cosh^{2} ax dx

∫ sinh^{2} ax cosh^{2} ax = \frac{sinh 4ax}{32a} − \frac{x}{8}

— * —

)clear all
--S 17 of 81
aa:=integrate(sinh(a*x)^2*cosh(a*x)^2,x)
--R
--R
--R                          3            3
--R        cosh(a x)sinh(a x)  + cosh(a x) sinh(a x) - a x
--R   (1)  -----------------------------------------------
--R                               8a
--R                                         Type: Union(Expression(Integer),...)
--E

--S 18 of 81
bb:=sinh(4*a*x)/(32*a)-x/8
--R
--R        sinh(4a x) - 4a x
--R   (2)  -----------------
--R               32a
--R                                                    Type: Expression(Integer)
--E
--S 19 of 81     14:594 Schaums and Axiom agree
cc:=complexNormalize(aa-bb)
--R
--R   (3)  0
--R                                                    Type: Expression(Integer)
--E

———-

6 [?]:14.595 ∫ \frac{dx}{sinh ax cosh ax}

∫ \frac{1}{sinh ax cosh ax} = \frac{1}{a} ln tanh ax

— * —

)clear all
--S 20 of 81
aa:=integrate(1/(sinh(a*x)*cosh(a*x)),x)
--R
--R
--R                      2cosh(a x)                     2sinh(a x)
--R        - log(- ---------------------) + log(- ---------------------)
--R                sinh(a x) - cosh(a x)          sinh(a x) - cosh(a x)
--R   (1)  -------------------------------------------------------------
--R                                      a
--R                                         Type: Union(Expression(Integer),...)
--E
--S 21 of 81
bb:=1/a*log(tanh(a*x))
--R
--R        log(tanh(a x))
--R   (2)  --------------
--R               a
--R                                                    Type: Expression(Integer)
--E

--S 22 of 81
cc:=aa-bb
--R
--R   (3)
--R                                      2cosh(a x)
--R       - log(tanh(a x)) - log(- ---------------------)
--R                                sinh(a x) - cosh(a x)
--R     +
--R                   2sinh(a x)
--R       log(- ---------------------)
--R             sinh(a x) - cosh(a x)
--R  /
--R     a
--R                                                    Type: Expression(Integer)
--E
--S 23 of 81
dd:=expandLog cc
--R
--R        - log(tanh(a x)) + log(sinh(a x)) - log(cosh(a x))
--R   (4)  --------------------------------------------------
--R                                 a
--R                                                    Type: Expression(Integer)
--E
--S 24 of 81
tanhrule:=rule(tanh(x) == sinh(x)/cosh(x))
--R
--R                   sinh(x)
--R   (5)  tanh(x) == -------
--R                   cosh(x)
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 25 of 81
ee:=tanhrule dd
--R
--R                             sinh(a x)
--R        log(sinh(a x)) - log(---------) - log(cosh(a x))
--R                             cosh(a x)
--R   (6)  ------------------------------------------------
--R                                a
--R                                                    Type: Expression(Integer)
--E
--S 26 of 81     14:595 Schaums and Axiom agree
ff:=expandLog ee
--R
--R   (7)  0
--R                                                    Type: Expression(Integer)

--E

———-

7 [?]:14.596 ∫ \frac{dx}{sinh^{2} ax cosh ax}

∫ \frac{1}{sinh^{2} ax cosh ax} = −\frac{1}{a} tan^{−1} sinh ax − \frac{csch ax}{a}

— * —

)clear all
--S 27 of 81
aa:=integrate(1/(sinh(a*x)^2*cosh(a*x)),x)
--R
--R   (1)
--R                      2                                   2
--R         (- 2sinh(a x)  - 4cosh(a x)sinh(a x) - 2cosh(a x)  + 2)
--R      *
--R         atan(sinh(a x) + cosh(a x))
--R     +
--R       - 2sinh(a x) - 2cosh(a x)
--R  /
--R                2                                      2
--R     a sinh(a x)  + 2a cosh(a x)sinh(a x) + a cosh(a x)  - a
--R                                         Type: Union(Expression(Integer),...)
--E
--S 28 of 81
bb:=-1/a*atan(sinh(a*x)-csch(a*x))/a
--R
--R          atan(sinh(a x) - csch(a x))
--R   (2)  - ---------------------------
--R                        2
--R                       a
--R                                                    Type: Expression(Integer)
--E
--S 29 of 81     14:596 Axiom cannot simplify this expression
cc:=aa-bb
--R
--R   (3)
--R                        2                                       2
--R         (- 2a sinh(a x)  - 4a cosh(a x)sinh(a x) - 2a cosh(a x)  + 2a)
--R      *
--R         atan(sinh(a x) + cosh(a x))
--R     +

--R                   2                                  2
--R         (sinh(a x)  + 2cosh(a x)sinh(a x) + cosh(a x)  - 1)
--R      *
--R         atan(sinh(a x) - csch(a x))
--R     +
--R       - 2a sinh(a x) - 2a cosh(a x)
--R  /
--R      2         2     2                      2         2    2
--R     a sinh(a x)  + 2a cosh(a x)sinh(a x) + a cosh(a x)  - a
--R                                                    Type: Expression(Integer)
--E

———-

8 [?]:14.597 ∫ \frac{dx}{sinh ax cosh^{2} ax}

∫ \frac{1}{sinh ax cosh^{2} ax} = \frac{sech ax}{a} + \frac{1}{a} ln tanh \frac{ax}{2}

— * —

)clear all
--S 30 of 81
aa:=integrate(1/(sinh(a*x)*cosh(a*x)^2),x)
--R
--R
--R   (1)
--R                     2                                  2
--R         (- sinh(a x)  - 2cosh(a x)sinh(a x) - cosh(a x)  - 1)
--R      *
--R         log(sinh(a x) + cosh(a x) + 1)
--R     +
--R                   2                                  2
--R         (sinh(a x)  + 2cosh(a x)sinh(a x) + cosh(a x)  + 1)
--R      *
--R         log(sinh(a x) + cosh(a x) - 1)
--R     +
--R       2sinh(a x) + 2cosh(a x)
--R  /
--R                2                                      2
--R     a sinh(a x)  + 2a cosh(a x)sinh(a x) + a cosh(a x)  + a
--R                                         Type: Union(Expression(Integer),...)
--E
--S 31 of 81
bb:=sech(a*x)/a+1/a*log(tanh((a*x)/2))
--R

--R                 a x
--R        log(tanh(---)) + sech(a x)
--R                  2
--R   (2)  --------------------------
--R                     a
--R                                                    Type: Expression(Integer)
--E
--S 32 of 81
cc:=aa-bb
--R
--R   (3)
--R                   2                                  2              a x
--R       (- sinh(a x)  - 2cosh(a x)sinh(a x) - cosh(a x)  - 1)log(tanh(---))
--R                                                                      2
--R     +
--R                     2                                  2
--R         (- sinh(a x)  - 2cosh(a x)sinh(a x) - cosh(a x)  - 1)
--R      *
--R         log(sinh(a x) + cosh(a x) + 1)
--R     +
--R                   2                                  2
--R         (sinh(a x)  + 2cosh(a x)sinh(a x) + cosh(a x)  + 1)
--R      *
--R         log(sinh(a x) + cosh(a x) - 1)
--R     +
--R                           2
--R       - sech(a x)sinh(a x)  + (- 2cosh(a x)sech(a x) + 2)sinh(a x)
--R     +
--R                   2
--R       (- cosh(a x)  - 1)sech(a x) + 2cosh(a x)
--R  /
--R                2                                      2
--R     a sinh(a x)  + 2a cosh(a x)sinh(a x) + a cosh(a x)  + a
--R                                                    Type: Expression(Integer)
--E
--S 33 of 81
sechrule:=rule(sech(x) == 1/cosh(x))
--R
--R                      1
--R   (4)  sech(x) == -------
--R                   cosh(x)
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 34 of 81
dd:=sechrule cc
--R
--R   (5)

--R                              2             2                     3
--R         (- cosh(a x)sinh(a x)  - 2cosh(a x) sinh(a x) - cosh(a x)  - cosh(a x))
--R      *
--R                  a x
--R         log(tanh(---))
--R                   2
--R     +
--R                              2             2                     3
--R         (- cosh(a x)sinh(a x)  - 2cosh(a x) sinh(a x) - cosh(a x)  - cosh(a x))
--R      *
--R         log(sinh(a x) + cosh(a x) + 1)
--R     +
--R                            2             2                     3
--R         (cosh(a x)sinh(a x)  + 2cosh(a x) sinh(a x) + cosh(a x)  + cosh(a x))
--R      *
--R         log(sinh(a x) + cosh(a x) - 1)
--R     +
--R                  2            2
--R       - sinh(a x)  + cosh(a x)  - 1
--R  /
--R                         2               2                       3
--R     a cosh(a x)sinh(a x)  + 2a cosh(a x) sinh(a x) + a cosh(a x)  + a cosh(a x)
--R                                                    Type: Expression(Integer)
--E
--S 35 of 81
tanhrule:=rule(tanh(x) == sinh(x)/cosh(x))
--R
--R                   sinh(x)
--R   (6)  tanh(x) == -------
--R                   cosh(x)
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 36 of 81
ee:=tanhrule dd
--R
--R   (7)
--R                              2             2                     3
--R         (- cosh(a x)sinh(a x)  - 2cosh(a x) sinh(a x) - cosh(a x)  - cosh(a x))
--R      *
--R         log(sinh(a x) + cosh(a x) + 1)
--R     +
--R                            2             2                     3
--R         (cosh(a x)sinh(a x)  + 2cosh(a x) sinh(a x) + cosh(a x)  + cosh(a x))
--R      *
--R         log(sinh(a x) + cosh(a x) - 1)
--R     +
--R                              2             2                     3
--R         (- cosh(a x)sinh(a x)  - 2cosh(a x) sinh(a x) - cosh(a x)  - cosh(a x))

--R      *
--R                  a x
--R             sinh(---)
--R                   2
--R         log(---------)
--R                  a x
--R             cosh(---)
--R                   2
--R     +
--R                  2            2
--R       - sinh(a x)  + cosh(a x)  - 1
--R  /
--R                         2               2                       3
--R     a cosh(a x)sinh(a x)  + 2a cosh(a x) sinh(a x) + a cosh(a x)  + a cosh(a x)
--R                                                    Type: Expression(Integer)
--E
--S 37 of 81
coshcuberule:=rule(cosh(x)^3 == 1/4*cosh(3*x)-3/4*cosh(x))
--R
--R               3    cosh(3x) - 3cosh(x)
--R   (8)  cosh(x)  == -------------------
--R                             4
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 38 of 81
ff:=coshcuberule ee
--R
--R   (9)
--R                                  2             2
--R             - 4cosh(a x)sinh(a x)  - 8cosh(a x) sinh(a x) - cosh(3a x)
--R           +
--R             - cosh(a x)
--R      *
--R         log(sinh(a x) + cosh(a x) + 1)
--R     +
--R                             2             2
--R         (4cosh(a x)sinh(a x)  + 8cosh(a x) sinh(a x) + cosh(3a x) + cosh(a x))
--R      *
--R         log(sinh(a x) + cosh(a x) - 1)
--R     +
--R                                  2             2
--R             - 4cosh(a x)sinh(a x)  - 8cosh(a x) sinh(a x) - cosh(3a x)
--R           +
--R             - cosh(a x)
--R      *
--R                  a x
--R             sinh(---)
--R                   2

--R         log(---------)
--R                  a x
--R             cosh(---)
--R                   2
--R     +
--R                   2             2
--R       - 4sinh(a x)  + 4cosh(a x)  - 4
--R  /
--R                            2               2
--R       4a cosh(a x)sinh(a x)  + 8a cosh(a x) sinh(a x) + a cosh(3a x)
--R     +
--R       a cosh(a x)
--R                                                    Type: Expression(Integer)
--E
--S 39 of 81
coshsqrrule:=rule(cosh(x)^2 == 1/2*cosh(2*x)+1/2)
--R
--R                2    cosh(2x) + 1
--R   (10)  cosh(x)  == ------------
--R                           2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 40 of 81
gg:=coshsqrrule ff
--R
--R   (11)
--R                                2
--R           - 4cosh(a x)sinh(a x)  + (- 4cosh(2a x) - 4)sinh(a x) - cosh(3a x)
--R         +
--R           - cosh(a x)
--R      *
--R         log(sinh(a x) + cosh(a x) + 1)
--R     +
--R                              2
--R           4cosh(a x)sinh(a x)  + (4cosh(2a x) + 4)sinh(a x) + cosh(3a x)
--R         +
--R           cosh(a x)
--R      *
--R         log(sinh(a x) + cosh(a x) - 1)
--R     +
--R                                2
--R           - 4cosh(a x)sinh(a x)  + (- 4cosh(2a x) - 4)sinh(a x) - cosh(3a x)
--R         +
--R           - cosh(a x)
--R      *
--R                  a x
--R             sinh(---)
--R                   2

--R         log(---------)
--R                  a x
--R             cosh(---)
--R                   2
--R     +
--R                   2
--R       - 4sinh(a x)  + 2cosh(2a x) - 2
--R  /
--R                            2
--R       4a cosh(a x)sinh(a x)  + (4a cosh(2a x) + 4a)sinh(a x) + a cosh(3a x)
--R     +
--R       a cosh(a x)
--R                                                    Type: Expression(Integer)
--E
--S 41 of 81
sinhsqrrule:=rule(sinh(x)^2 == 1/2*cosh(2*x)-1/2)
--R
--R                2    cosh(2x) - 1
--R   (12)  sinh(x)  == ------------
--R                           2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 42 of 81
hh:=sinhsqrrule gg
--R
--R   (13)
--R       - log(sinh(a x) + cosh(a x) + 1) + log(sinh(a x) + cosh(a x) - 1)
--R     +
--R                  a x
--R             sinh(---)
--R                   2
--R       - log(---------)
--R                  a x
--R             cosh(---)
--R                   2
--R  /
--R     a
--R                                                    Type: Expression(Integer)
--E
--S 43 of 81
ii:=expandLog hh
--R
--R   (14)
--R       - log(sinh(a x) + cosh(a x) + 1) + log(sinh(a x) + cosh(a x) - 1)
--R     +
--R                  a x              a x
--R       - log(sinh(---)) + log(cosh(---))

--R                   2                2
--R  /
--R     a
--R                                                    Type: Expression(Integer)
--E
--S 44 of 81     14:597 Schaums and Axiom agree
jj:=complexNormalize ii
--R
--R   (15)  0
--R                                                    Type: Expression(Integer)
--E

———-

9 [?]:14.598 ∫ \frac{dx}{sinh^{2} ax cosh^{2} ax}

∫ \frac{1}{sinh^{2} ax cosh^{2} ax} = −\frac{2 coth 2ax}{a}

— * —

)clear all
--S 45 of 81
aa:=integrate(1/(sinh(a*x)^2*cosh(a*x)^2),x)
--R
--R
--R   (1)
--R   -
--R        4
--R     /
--R                     4                        3               2         2
--R          a sinh(a x)  + 4a cosh(a x)sinh(a x)  + 6a cosh(a x) sinh(a x)
--R        +
--R                      3                       4
--R          4a cosh(a x) sinh(a x) + a cosh(a x)  - a
--R                                         Type: Union(Expression(Integer),...)
--E
--S 46 of 81
bb:=-(2*coth(2*a*x))/a
--R
--R          2coth(2a x)
--R   (2)  - -----------
--R               a
--R                                                    Type: Expression(Integer)
--E

--S 47 of 81     14:598 Axiom cannot simplify this expression
cc:=aa-bb
--R
--R   (3)
--R                           4                                3
--R       2coth(2a x)sinh(a x)  + 8cosh(a x)coth(2a x)sinh(a x)
--R     +
--R                  2                   2             3
--R       12cosh(a x) coth(2a x)sinh(a x)  + 8cosh(a x) coth(2a x)sinh(a x)
--R     +
--R                  4
--R       (2cosh(a x)  - 2)coth(2a x) - 4
--R  /
--R                  4                        3               2         2
--R       a sinh(a x)  + 4a cosh(a x)sinh(a x)  + 6a cosh(a x) sinh(a x)
--R     +
--R                   3                       4
--R       4a cosh(a x) sinh(a x) + a cosh(a x)  - a
--R                                                    Type: Expression(Integer)
--E

———-

10 [?]:14.599 ∫ \begin{matrix}sinh & 2 & ax \\ cosh & ax \end{matrix} dx

∫ \begin{matrix}sinh & 2 & ax \\ cosh & ax \end{matrix} dx = \frac{sinh ax}{a} − \frac{1}{a} tan^{−1} sinh ax

— * —

)clear all
--S 48 of 81
aa:=integrate(sinh(a*x)^2/cosh(a*x),x)
--R
--R
--R   (1)
--R                                                                         2
--R       (- 4sinh(a x) - 4cosh(a x))atan(sinh(a x) + cosh(a x)) + sinh(a x)
--R     +
--R                                      2
--R       2cosh(a x)sinh(a x) + cosh(a x)  - 1
--R  /
--R     2a sinh(a x) + 2a cosh(a x)
--R                                         Type: Union(Expression(Integer),...)
--E

--S 49 of 81
bb:=sinh(a*x)/a-1/a*atan(sinh(a*x))
--R
--R        - atan(sinh(a x)) + sinh(a x)
--R   (2)  -----------------------------
--R                      a
--R                                                    Type: Expression(Integer)
--E
--S 50 of 81     14:599 Axiom cannot simplify this expression
cc:=aa-bb
--R
--R   (3)
--R       (- 4sinh(a x) - 4cosh(a x))atan(sinh(a x) + cosh(a x))
--R     +
--R                                                           2            2
--R       (2sinh(a x) + 2cosh(a x))atan(sinh(a x)) - sinh(a x)  + cosh(a x)  - 1
--R  /
--R     2a sinh(a x) + 2a cosh(a x)
--R                                                    Type: Expression(Integer)
--E

———-

11 [?]:14.600 ∫ \begin{matrix}cosh & 2 & ax \\ sinh & ax \end{matrix} dx

∫ \begin{matrix}cosh & 2 & ax \\ sinh & ax \end{matrix} = \frac{cosh ax}{a} + \frac{1}{a} ln tanh \frac{ax}{2}

— * —

)clear all
--S 51 of 81
aa:=integrate(cosh(a*x)^2/sinh(a*x),x)
--R
--R
--R   (1)
--R       (- 2sinh(a x) - 2cosh(a x))log(sinh(a x) + cosh(a x) + 1)
--R     +
--R                                                                          2
--R       (2sinh(a x) + 2cosh(a x))log(sinh(a x) + cosh(a x) - 1) + sinh(a x)
--R     +
--R                                      2
--R       2cosh(a x)sinh(a x) + cosh(a x)  + 1
--R  /
--R     2a sinh(a x) + 2a cosh(a x)
--R                                         Type: Union(Expression(Integer),...)

--E
--S 52 of 81
bb:=cosh(a*x)/a+1/a*log(tanh((a*x)/2))
--R
--R                 a x
--R        log(tanh(---)) + cosh(a x)
--R                  2
--R   (2)  --------------------------
--R                     a
--R                                                    Type: Expression(Integer)
--E
--S 53 of 81     14:600 Axiom cannot simplify this expression
cc:=aa-bb
--R
--R   (3)
--R                                           a x
--R       (- 2sinh(a x) - 2cosh(a x))log(tanh(---))
--R                                            2
--R     +
--R       (- 2sinh(a x) - 2cosh(a x))log(sinh(a x) + cosh(a x) + 1)
--R     +
--R                                                                          2
--R       (2sinh(a x) + 2cosh(a x))log(sinh(a x) + cosh(a x) - 1) + sinh(a x)
--R     +
--R                  2
--R       - cosh(a x)  + 1
--R  /
--R     2a sinh(a x) + 2a cosh(a x)
--R                                                    Type: Expression(Integer)
--E

———-

12 [?]:14.601 ∫ \frac{dx}{cosh ax(1 + sinh ax)}

∫ \frac{1}{cosh ax(1 + sinh ax)} = \frac{1}{2a} ln (\frac{1 + sinh ax}{cosh ax}) + \frac{1}{a} tan^{−1} e^{ax}

— * —

)clear all
--S 54 of 81
aa:=integrate(1/(cosh(a*x)*(1+sinh(a*x))),x)
--R
--R

--R   (1)
--R                     2cosh(a x)                - 2sinh(a x) - 2
--R       - log(- ---------------------) + log(---------------------)
--R               sinh(a x) - cosh(a x)        sinh(a x) - cosh(a x)
--R     +
--R       2atan(sinh(a x) + cosh(a x))
--R  /
--R     2a
--R                                         Type: Union(Expression(Integer),...)
--E
--S 55 of 81
bb:=1/(2*a)*log((1+sinh(a*x))/cosh(a*x))+1/a*atan(%e^(a*x))
--R
--R            sinh(a x) + 1            a x
--R        log(-------------) + 2atan(%e   )
--R              cosh(a x)
--R   (2)  ---------------------------------
--R                        2a
--R                                                    Type: Expression(Integer)
--E
--S 56 of 81
cc:=aa-bb
--R
--R   (3)
--R             sinh(a x) + 1                2cosh(a x)
--R       - log(-------------) - log(- ---------------------)
--R               cosh(a x)            sinh(a x) - cosh(a x)
--R     +
--R              - 2sinh(a x) - 2                                             a x
--R       log(---------------------) + 2atan(sinh(a x) + cosh(a x)) - 2atan(%e   )
--R           sinh(a x) - cosh(a x)
--R  /
--R     2a
--R                                                    Type: Expression(Integer)
--E
--S 57 of 81
dd:=expandLog cc
--R
--R                                             a x
--R        atan(sinh(a x) + cosh(a x)) - atan(%e   )
--R   (4)  -----------------------------------------
--R                            a
--R                                                    Type: Expression(Integer)
--E
--S 58 of 81
atanrule:=rule(atan(x) == -%i/2*log((1+%i*x)/(1-%i*x)))

--R
--R                            - x + %i
--R                     %i log(--------)
--R                             x + %i
--R   (5)  atan(x) == - ----------------
--R                             2
--R     Type: RewriteRule(Integer,Complex(Integer),Expression(Complex(Integer)))
--E
--S 59 of 81
ee:=atanrule dd
--R
--R                   a x
--R               - %e    + %i           - sinh(a x) - cosh(a x) + %i
--R        %i log(------------) - %i log(----------------------------)
--R                  a x                  sinh(a x) + cosh(a x) + %i
--R                %e    + %i
--R   (6)  -----------------------------------------------------------
--R                                     2a
--R                                           Type: Expression(Complex(Integer))
--E
--S 60 of 81
ff:=expandLog ee
--R
--R   (7)
--R       %i log(sinh(a x) + cosh(a x) + %i) - %i log(sinh(a x) + cosh(a x) - %i)
--R     +
--R                  a x                  a x
--R       - %i log(%e    + %i) + %i log(%e    - %i)
--R  /
--R     2a
--R                                           Type: Expression(Complex(Integer))
--E
--S 61 of 81     14:601 Schaums and Axiom agree
gg:=complexNormalize ff
--R
--R   (8)  0
--R                                           Type: Expression(Complex(Integer))
--E

———-

13 [?]:14.602 ∫ \frac{dx}{sinh ax(cosh ax + 1)}

∫ \frac{1}{sinh ax(cosh ax + 1)} = \frac{1}{2a} ln tanh \frac{ax}{2} + \frac{1}{2a(cosh ax + 1)}

— * —

)clear all
--S 62 of 81
aa:=integrate(1/(sinh(a*x)*(cosh(a*x)+1)),x)
--R
--R
--R   (1)
--R                      2                                          2
--R           - sinh(a x)  + (- 2cosh(a x) - 2)sinh(a x) - cosh(a x)  - 2cosh(a x)
--R         +
--R           - 1
--R      *
--R         log(sinh(a x) + cosh(a x) + 1)
--R     +
--R                   2                                        2
--R         (sinh(a x)  + (2cosh(a x) + 2)sinh(a x) + cosh(a x)  + 2cosh(a x) + 1)
--R      *
--R         log(sinh(a x) + cosh(a x) - 1)
--R     +
--R       2sinh(a x) + 2cosh(a x)
--R  /
--R                   2                                              2
--R       2a sinh(a x)  + (4a cosh(a x) + 4a)sinh(a x) + 2a cosh(a x)
--R     +
--R       4a cosh(a x) + 2a
--R                                         Type: Union(Expression(Integer),...)
--E
--S 63 of 81
bb:=1/(2*a)*log(tanh((a*x)/2))+1/(2*a*(cosh(a*x)+1))
--R
--R                                a x
--R        (cosh(a x) + 1)log(tanh(---)) + 1
--R                                 2
--R   (2)  ---------------------------------
--R                2a cosh(a x) + 2a
--R                                                    Type: Expression(Integer)
--E
--S 64 of 81
cc:=aa-bb
--R
--R   (3)
--R                                     2
--R           (- cosh(a x) - 1)sinh(a x)
--R         +
--R                        2                                       3             2

--R           (- 2cosh(a x)  - 4cosh(a x) - 2)sinh(a x) - cosh(a x)  - 3cosh(a x)
--R         +
--R           - 3cosh(a x) - 1
--R      *
--R                  a x
--R         log(tanh(---))
--R                   2
--R     +
--R                                     2
--R           (- cosh(a x) - 1)sinh(a x)
--R         +
--R                        2                                       3             2
--R           (- 2cosh(a x)  - 4cosh(a x) - 2)sinh(a x) - cosh(a x)  - 3cosh(a x)
--R         +
--R           - 3cosh(a x) - 1
--R      *
--R         log(sinh(a x) + cosh(a x) + 1)
--R     +
--R                                   2              2
--R           (cosh(a x) + 1)sinh(a x)  + (2cosh(a x)  + 4cosh(a x) + 2)sinh(a x)
--R         +
--R                    3             2
--R           cosh(a x)  + 3cosh(a x)  + 3cosh(a x) + 1
--R      *
--R         log(sinh(a x) + cosh(a x) - 1)
--R     +
--R                  2            2
--R       - sinh(a x)  + cosh(a x)  - 1
--R  /
--R                                   2
--R       (2a cosh(a x) + 2a)sinh(a x)
--R     +
--R                    2                                             3
--R       (4a cosh(a x)  + 8a cosh(a x) + 4a)sinh(a x) + 2a cosh(a x)
--R     +
--R                   2
--R       6a cosh(a x)  + 6a cosh(a x) + 2a
--R                                                    Type: Expression(Integer)
--E
--S 65 of 81
coshcuberule:=rule(cosh(x)^3 == 1/4*cosh(3*x)-3/4*cosh(x))
--R
--R               3    cosh(3x) - 3cosh(x)
--R   (4)  cosh(x)  == -------------------
--R                             4
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 66 of 81

dd:=coshcuberule cc
--R
--R   (5)
--R                                      2
--R           (- 4cosh(a x) - 4)sinh(a x)
--R         +
--R                        2
--R           (- 8cosh(a x)  - 16cosh(a x) - 8)sinh(a x) - cosh(3a x)
--R         +
--R                        2
--R           - 12cosh(a x)  - 9cosh(a x) - 4
--R      *
--R                  a x
--R         log(tanh(---))
--R                   2
--R     +
--R                                      2
--R           (- 4cosh(a x) - 4)sinh(a x)
--R         +
--R                        2
--R           (- 8cosh(a x)  - 16cosh(a x) - 8)sinh(a x) - cosh(3a x)
--R         +
--R                        2
--R           - 12cosh(a x)  - 9cosh(a x) - 4
--R      *
--R         log(sinh(a x) + cosh(a x) + 1)
--R     +
--R                                    2              2
--R           (4cosh(a x) + 4)sinh(a x)  + (8cosh(a x)  + 16cosh(a x) + 8)sinh(a x)
--R         +
--R                                   2
--R           cosh(3a x) + 12cosh(a x)  + 9cosh(a x) + 4
--R      *
--R         log(sinh(a x) + cosh(a x) - 1)
--R     +
--R                   2             2
--R       - 4sinh(a x)  + 4cosh(a x)  - 4
--R  /
--R                                   2
--R       (8a cosh(a x) + 8a)sinh(a x)
--R     +
--R                     2
--R       (16a cosh(a x)  + 32a cosh(a x) + 16a)sinh(a x) + 2a cosh(3a x)
--R     +
--R                    2
--R       24a cosh(a x)  + 18a cosh(a x) + 8a
--R                                                    Type: Expression(Integer)
--E
--S 67 of 81

sinhsqrrule:=rule(sinh(x)^2 == 1/2*cosh(2*x)-1/2)
--R
--R               2    cosh(2x) - 1
--R   (6)  sinh(x)  == ------------
--R                          2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 68 of 81
ee:=sinhsqrrule dd
--R
--R   (7)
--R                        2
--R           (- 8cosh(a x)  - 16cosh(a x) - 8)sinh(a x) - cosh(3a x)
--R         +
--R                                                     2
--R           (- 2cosh(a x) - 2)cosh(2a x) - 12cosh(a x)  - 7cosh(a x) - 2
--R      *
--R                  a x
--R         log(tanh(---))
--R                   2
--R     +
--R                        2
--R           (- 8cosh(a x)  - 16cosh(a x) - 8)sinh(a x) - cosh(3a x)
--R         +
--R                                                     2
--R           (- 2cosh(a x) - 2)cosh(2a x) - 12cosh(a x)  - 7cosh(a x) - 2
--R      *
--R         log(sinh(a x) + cosh(a x) + 1)
--R     +
--R                      2
--R           (8cosh(a x)  + 16cosh(a x) + 8)sinh(a x) + cosh(3a x)
--R         +
--R                                                   2
--R           (2cosh(a x) + 2)cosh(2a x) + 12cosh(a x)  + 7cosh(a x) + 2
--R      *
--R         log(sinh(a x) + cosh(a x) - 1)
--R     +
--R                                 2
--R       - 2cosh(2a x) + 4cosh(a x)  - 2
--R  /
--R                     2
--R       (16a cosh(a x)  + 32a cosh(a x) + 16a)sinh(a x) + 2a cosh(3a x)
--R     +
--R                                                    2
--R       (4a cosh(a x) + 4a)cosh(2a x) + 24a cosh(a x)  + 14a cosh(a x) + 4a
--R                                                    Type: Expression(Integer)
--E
--S 69 of 81

coshsqrrule:=rule(cosh(x)^2 == 1/2*cosh(2*x)+1/2)
--R
--R               2    cosh(2x) + 1
--R   (8)  cosh(x)  == ------------
--R                          2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 70 of 81
ff:=coshsqrrule ee
--R
--R   (9)
--R                  a x
--R       - log(tanh(---)) - log(sinh(a x) + cosh(a x) + 1)
--R                   2
--R     +
--R       log(sinh(a x) + cosh(a x) - 1)
--R  /
--R     2a
--R                                                    Type: Expression(Integer)
--E
--S 71 of 81     14:602 Schaums and Axiom agree
gg:=complexNormalize ff
--R
--R   (10)  0
--R                                                    Type: Expression(Integer)
--E

———-

14 [?]:14.603 ∫ \frac{dx}{sinh ax(cosh ax − 1)}

∫ \frac{1}{sinh ax(cosh ax − 1)} = −\frac{1}{2a} ln tanh \frac{ax}{2} − \frac{1}{2a(coshax − 1)}

— * —

)clear all
--S 72 of 81
aa:=integrate(1/(sinh(a*x)*(cosh(a*x)-1)),x)
--R
--R
--R   (1)
--R                   2                                        2
--R         (sinh(a x)  + (2cosh(a x) - 2)sinh(a x) + cosh(a x)  - 2cosh(a x) + 1)
--R      *

--R         log(sinh(a x) + cosh(a x) + 1)
--R     +
--R                      2                                          2
--R           - sinh(a x)  + (- 2cosh(a x) + 2)sinh(a x) - cosh(a x)  + 2cosh(a x)
--R         +
--R           - 1
--R      *
--R         log(sinh(a x) + cosh(a x) - 1)
--R     +
--R       - 2sinh(a x) - 2cosh(a x)
--R  /
--R                   2                                              2
--R       2a sinh(a x)  + (4a cosh(a x) - 4a)sinh(a x) + 2a cosh(a x)
--R     +
--R       - 4a cosh(a x) + 2a
--R                                         Type: Union(Expression(Integer),...)
--E
--S 73 of 81
bb:=-1/(2*a)*log(tanh((a*x)/2))-1/(2*a*(cosh(a*x)-1))
--R
--R                                  a x
--R        (- cosh(a x) + 1)log(tanh(---)) - 1
--R                                   2
--R   (2)  -----------------------------------
--R                 2a cosh(a x) - 2a
--R                                                    Type: Expression(Integer)
--E
--S 74 of 81
cc:=aa-bb
--R
--R   (3)
--R                                   2              2
--R           (cosh(a x) - 1)sinh(a x)  + (2cosh(a x)  - 4cosh(a x) + 2)sinh(a x)
--R         +
--R                    3             2
--R           cosh(a x)  - 3cosh(a x)  + 3cosh(a x) - 1
--R      *
--R                  a x
--R         log(tanh(---))
--R                   2
--R     +
--R                                   2              2
--R           (cosh(a x) - 1)sinh(a x)  + (2cosh(a x)  - 4cosh(a x) + 2)sinh(a x)
--R         +
--R                    3             2
--R           cosh(a x)  - 3cosh(a x)  + 3cosh(a x) - 1
--R      *
--R         log(sinh(a x) + cosh(a x) + 1)

--R     +
--R                                     2
--R           (- cosh(a x) + 1)sinh(a x)
--R         +
--R                        2                                       3             2
--R           (- 2cosh(a x)  + 4cosh(a x) - 2)sinh(a x) - cosh(a x)  + 3cosh(a x)
--R         +
--R           - 3cosh(a x) + 1
--R      *
--R         log(sinh(a x) + cosh(a x) - 1)
--R     +
--R                2            2
--R       sinh(a x)  - cosh(a x)  + 1
--R  /
--R                                   2
--R       (2a cosh(a x) - 2a)sinh(a x)
--R     +
--R                    2                                             3
--R       (4a cosh(a x)  - 8a cosh(a x) + 4a)sinh(a x) + 2a cosh(a x)
--R     +
--R                     2
--R       - 6a cosh(a x)  + 6a cosh(a x) - 2a
--R                                                    Type: Expression(Integer)
--E
--S 75 of 81
coshcuberule:=rule(cosh(x)^3 == 1/4*cosh(3*x)-3/4*cosh(x))
--R
--R               3    cosh(3x) - 3cosh(x)
--R   (4)  cosh(x)  == -------------------
--R                             4
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 76 of 81
dd:=coshcuberule cc
--R
--R   (5)
--R                                    2              2
--R           (4cosh(a x) - 4)sinh(a x)  + (8cosh(a x)  - 16cosh(a x) + 8)sinh(a x)
--R         +
--R                                   2
--R           cosh(3a x) - 12cosh(a x)  + 9cosh(a x) - 4
--R      *
--R                  a x
--R         log(tanh(---))
--R                   2
--R     +
--R                                    2              2
--R           (4cosh(a x) - 4)sinh(a x)  + (8cosh(a x)  - 16cosh(a x) + 8)sinh(a x)

--R         +
--R                                   2
--R           cosh(3a x) - 12cosh(a x)  + 9cosh(a x) - 4
--R      *
--R         log(sinh(a x) + cosh(a x) + 1)
--R     +
--R                                      2
--R           (- 4cosh(a x) + 4)sinh(a x)
--R         +
--R                        2
--R           (- 8cosh(a x)  + 16cosh(a x) - 8)sinh(a x) - cosh(3a x)
--R         +
--R                      2
--R           12cosh(a x)  - 9cosh(a x) + 4
--R      *
--R         log(sinh(a x) + cosh(a x) - 1)
--R     +
--R                 2             2
--R       4sinh(a x)  - 4cosh(a x)  + 4
--R  /
--R                                   2
--R       (8a cosh(a x) - 8a)sinh(a x)
--R     +
--R                     2
--R       (16a cosh(a x)  - 32a cosh(a x) + 16a)sinh(a x) + 2a cosh(3a x)
--R     +
--R                      2
--R       - 24a cosh(a x)  + 18a cosh(a x) - 8a
--R                                                    Type: Expression(Integer)
--E
--S 77 of 81
sinhsqrrule:=rule(sinh(x)^2 == 1/2*cosh(2*x)-1/2)
--R
--R               2    cosh(2x) - 1
--R   (6)  sinh(x)  == ------------
--R                          2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 78 of 81
ee:=sinhsqrrule dd
--R
--R   (7)
--R                      2
--R           (8cosh(a x)  - 16cosh(a x) + 8)sinh(a x) + cosh(3a x)
--R         +
--R                                                   2
--R           (2cosh(a x) - 2)cosh(2a x) - 12cosh(a x)  + 7cosh(a x) - 2
--R      *

--R                  a x
--R         log(tanh(---))
--R                   2
--R     +
--R                      2
--R           (8cosh(a x)  - 16cosh(a x) + 8)sinh(a x) + cosh(3a x)
--R         +
--R                                                   2
--R           (2cosh(a x) - 2)cosh(2a x) - 12cosh(a x)  + 7cosh(a x) - 2
--R      *
--R         log(sinh(a x) + cosh(a x) + 1)
--R     +
--R                        2
--R           (- 8cosh(a x)  + 16cosh(a x) - 8)sinh(a x) - cosh(3a x)
--R         +
--R                                                     2
--R           (- 2cosh(a x) + 2)cosh(2a x) + 12cosh(a x)  - 7cosh(a x) + 2
--R      *
--R         log(sinh(a x) + cosh(a x) - 1)
--R     +
--R                               2
--R       2cosh(2a x) - 4cosh(a x)  + 2
--R  /
--R                     2
--R       (16a cosh(a x)  - 32a cosh(a x) + 16a)sinh(a x) + 2a cosh(3a x)
--R     +
--R                                                    2
--R       (4a cosh(a x) - 4a)cosh(2a x) - 24a cosh(a x)  + 14a cosh(a x) - 4a
--R                                                    Type: Expression(Integer)
--E
--S 79 of 81
coshsqrrule:=rule(cosh(x)^2 == 1/2*cosh(2*x)+1/2)
--R
--R               2    cosh(2x) + 1
--R   (8)  cosh(x)  == ------------
--R                          2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 80 of 81
ff:=coshsqrrule ee
--R
--R   (9)
--R                a x
--R       log(tanh(---)) + log(sinh(a x) + cosh(a x) + 1)
--R                 2
--R     +
--R       - log(sinh(a x) + cosh(a x) - 1)
--R  /

--R     2a
--R                                                    Type: Expression(Integer)
--E
--S 81 of 81     14:603 Schaums and Axiom agree
gg:=complexNormalize ff
--R
--R   (10)  0
--R                                                    Type: Expression(Integer)
--E
)spool
)lisp (bye)

———-

References

[1] Spiegel, Murray R. Mathematical Handbook of Formulas and Tables Schaum’s Outline Series McGraw-Hill 1968 pp89-90